# Large Language Models Could Be Rote Learners

### v4

@@ Appendix B Details of TrinEval
Prompt template for triplet extraction:
You are an expert of Knowledge Keyword extraction. Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question. Given a Fact corpus, a Question about the Fact corpus, and the Answer to the Question, analyze the Question corpus as well as the given Answer. Applying the provided steps, extract the Knowledge Keyword, the Attribute of the Knowledge Keyword and the necessary Context to obtain the key information of the Question, ensuring they are sufficient for answering the given Question and obtaining the given Answer. # Steps 1. **Review the Fact corpus:** Read through the entire Fact corpus to understand the context. 2. **Identify the Question:** Focus on the given Question to capture which part of the Fact corpus it is asking about. 3. **Understand the Answer to the Question:** Compare the given Answer and the identified questioned part within the Fact corpus and understand why this answer was chosen. 4. **Write Step-by-Step Reasoning:** - Identify the asked Knowledge Keyword in the Question that is the subject of the most information in the Fact corpus and the asked Question is about the information among. - Determine the asked Attribute of the Knowledge Keyword in the Question, which can be used to infer the given Answer. - Review the identified Knowledge Keyword and Attribute to confirm that only these two parts can be used to obtain the given Answer to the given Question. If not, extract all the necessary Context from the Question that makes it enough to obtain the given Answer to the given Question. 5. **Determine Outcome:** Based on the reasoning, conclude and extract the Knowledge Keyword, the Attribute and the Context (if necessary) of the Question according to the Question corpus. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Knowledge Keyword:** [Extracted Knowledge Keyword here] - **Attribute:** [Extracted Attribute of the Knowledge Keyword here] - **Context:** [Extracted Context within the Question to make up for the Knowledge Keyword and the Attribute here if necessary] # Examples [examples] # Notes - Strictly follow the format of the examples and give Knowledge Keywords, the Attribute and the Context (if necessary) anyway. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should be the original text within the Question and should not incorporate any phrases that cannot be exactly matched in the Question. - Never include any information from the options of the multiple choice question, especially the content of the answer option. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should include all the necessary information only within the Question Corpus for answering the Question and obtaining the given Answer. **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
Prompt template for triplet validation & reflection:
You are an expert of [subject] and an advanced reasoning agent that can determine whether the given Knowledge Keyword, Attribute of the Knowledge Keyword and the Context present most of the necessary information of the Question for obtaining the given Answer. Suppose you have sufficient background knowledge about subj. Consider the given Knowledge Keyword, Attribute and the Context, then determine whether the given Answer can be directly obtained from them even without the Question. # Steps 1. **Check the Semantic completeness:** Suppose you have sufficient background knowledge about [subject], and you can solve the given Question and obtain the given Answer. Read through the given Knowledge Keyword, Attribute, Context and the given Question. Check if the given Knowledge Keyword, Attribute, Context are the original text within the Question and contain the necessary queried information the Question itself provided (ignore the information the Question did not provided). If not so, check if the missed information is indeed incorporated in the Question (which is not acceptable, but if not, it is acceptable). Point out the information that is within the Question but they have missed. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. 2. **Check the Answer relevance:** Suppose you have sufficient background knowledge about subj, and you can solve the given Question and obtain the given Answer. Read through the given Knowledge Keyword, Attribute, Context and the given Question. Read through the given Knowledge Keyword, Attribute, Context and the given Answer. Check if the Answer can be directly inferred with the given Knowledge Keyword, Attribute and the Context without seeing the Question. If not so, check if the missed information is indeed incorporated in the Question (which is not acceptable, but if not, it is acceptable). Point out the information that is within the Question but they have missed. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. 3. **Check the Semantic Redundancy:** Read through the given Knowledge Keyword, Attribute, Context, the given Question and the given corresponding Answer. Check if the Answer can be directly matched within the given Knowledge Keyword, Attribute and the Context. Check if there are any unnecessary information within the given Knowledge Keyword, Attribute and the Context for obtaining the given Answer to the Question. If not so, point out what is redundant. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Verdict for the given Knowledge Keyword, Attribute and Context:** [Single verdict (Yes/No) here for whether the given Knowledge Keyword, Attribute and Context contain most of the asked information of the Question, can be used to infer the given Answer with only them without the whole Question, and do not contain redundant information for obtaining the given Answer.] # Notes - Do not deviate from the specified format. Do not generate anything else after the Verdict (only Yes/No) for the given Knowledge Keyword, Attribute and Context. - Suppose you have sufficient background knowledge about subj, and you can solve the given Question and obtain the given Answer. For Semantic completeness and Answer relevance, it is acceptable to miss information that is also not incorporated in the Question. - Provide a detailed explanation following the given steps before arriving at the verdict (Yes/No). Provide a final verdict (only Yes/No) in order at the end in the given format. - **Question:** [question] - **Answer:** [answer] - **Knowledge Keyword:** [extracted knowledge entity] - **Attribute:** [extracted attribute] - **Context:** [extracted context]
Prompt template for the second round triplet extraction:
You are an advanced reasoning agent that can improve through self-reflection and an expert of Knowledge Keyword extraction. Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question. Given a Fact corpus, a Question about the Fact corpus, and the Answer to the Question, analyze the Question corpus as well as the given Answer. Applying the provided steps, extract the Knowledge Keyword, the Attribute of the Knowledge Keyword and the necessary Context to rephrase the Question, ensuring they are sufficient for answering the given Question and obtaining the given Answer. # Steps 1. **Review the Fact corpus:** Read through the entire Fact corpus to understand the context. 2. **Identify the Question:** Focus on the given Question to capture which part of the Fact corpus it is asking about. 3. **Understand the Answer to the Question:** Compare the given Answer and the identified questioned part within the Fact corpus and understand why this answer was chosen. 4. **Write Step-by-Step Reasoning:** - Identify the asked Knowledge Keyword in the Question that is the subject of the most information in the Fact corpus and the asked Question is about the information among. - Determine the asked Attribute of the Knowledge Keyword in the Question, which can be used to infer the given Answer. - Review the identified Knowledge Keyword and Attribute to confirm that only these two parts can be used to obtain the given Answer to the given Question. If not, extract all the necessary Context from the Question that makes it enough to obtain the given Answer to the given Question. 5. **Determine Outcome:** Based on the reasoning, conclude and extract the Knowledge Keyword, the Attribute and the Context (if necessary) of the Question according to the Question corpus. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Knowledge Keyword:** [Extracted Knowledge Keyword here] - **Attribute:** [Extracted Attribute of the Knowledge Keyword here] - **Context:** [Extracted Context within the Question to make up for the Knowledge Keyword and the Attribute here if necessary] # Examples [examples] You will be given a previous trial. You were unsuccessful in extracting the Knowledge Keyword, Attribute and the necessary that meet the requirements in the previous trial. Given the Reflection below, improve the process. The process is as follows: # Previous returns: - **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] - **Question:** [question] - **Answer:** [answer option content] - **Knowledge Keyword:** [extracted knowledge entity of the last trial] - **Attribute:** [attribute of the last trial] - **Context:** [context of the last trial] - **Reflection:** [rational of the last trial] # Notes - Consider the Reflection given above. Improve the extraction of Knowledge Keyword, Attribute and Context (if necessary). - Strictly follow the format of the examples and give Knowledge Keywords, the Attribute and the Context (if necessary) anyway. - The extracted Knowledge Keyword should be phrases within the Question and should not incorporate any information of the Fact corpus or the given Answer that is not mentioned in the Question. - The extracted Attribute and Context (if necessary) should only include information from the Question corpus. Never include information from the options of the multiple choice question, especially the content of the answer option. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should include all the necessary information only within the Question Corpus for answering the Question and obtaining the given Answer. **Fact:** [question] [option content list] [subject] [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
Original MCQ Original MCQ Example
You are an expert on multiple choice questions of [subject]. Analyze the given question and the given options. Determine the correct answer option to the question. Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options. Generate the option ID of the correct option (answer). - **Question:** [question] - **Options:** A. [option A] B. [option B] C. [option C] D. [option D] You are an expert on multiple choice questions of high school computer science. Analyze the given question and the given options. Determine the correct answer option to the question. Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options. Generate the option ID of the correct option (answer). - **Question:** Which of the following is usually NOT represented in a subroutine’s activation record frame for a stack-based programming language? - **Options:** A. Values of local variables B. A heap area C. The return address D. Stack pointer for the calling activation record
TrinEval MCQ TrinEval MCQ Example
[…]

### v5

@@ III-B Measuring LLM Capability with TrinEval
Finally, prompting with the extracted $x_{E}$, $x_{A}$, $x_{C}$, and options $x_{O}$ (see Fig. 5 in Appendix B for illustration), we inspect the generation probability of the next ground-truth answer token $x_{W}$ (i.e., A/B/C/D) as the measurement:
$F_{c}(x,G)=p_{G}(x_{W}|x_{E},x_{A},x_{C},x_{O}).$ (1)
As can be seen, $F_{c}$ metric retains the necessary knowledge-centric information while discarding unnecessary biases, especially the rote memorization of LLMs, which leads to the quantification of genuine capability of LLMs.
III-C Quantifying LLM Memorization
On the other hand, in order to look into the nuanced forms of learning, we believe that a continuous metric is required for memorization evaluation. For quantifying the memorization of LLMs, prior research [16, 17] suggests that outlier tokens, which exhibit higher generation probabilities [^5] As established in Shi et al. [16]: “an unseen example is likely to contain a few outlier words with low probabilities under the LLM, while a seen example is less likely to have words with such low probabilities.”, are more likely to be found in memorized samples. Experiments on WIKIMIA [16] also finds that this method exhibits high AUC score on seen samples. In order to avoid discrete binary classification, building on this idea, we adopt the metric that utilizes the bottom $K\%$ of token probabilities within the generated sequence as a measure of memorization (we primarily set the value of $K$ to 10 in this paper). Formally, the memorization score $F_{m}(\overline{x},G)$ of LLM $G$ on text sequence $\overline{x}$ is computed as follows [^6] In the experiment, we have also tested other similar metrics like Min-K++% [17], which show consistent results (exhibit a Pearson correlation coefficient of 0.9007 (strong positive correlation)). Thus, we only use Min-K% [16] as an example here:
where $p_{G}(\overline{x}_{i}|\overline{x}_{1:i-1})$ denotes the generation probability of token $\overline{x}_{i}$ by $G$ given its prefix subsequence as context, and set $\mathcal{M}_{K}(\overline{x})$ includes the $K\%$ of tokens with the lowest probabilities. The higher $F_{m}$ is, the more likely $\overline{x}$ is memorized by the LLM, i.e., the least memorized content could still been extracted with a high probability. Note that our objective is not to achieve high-precision quantification of memorization. Instead, we aim to statistically demonstrate a valid quantification of memorization across vast samples for comparative analysis.
IV Experiments
Q3. What insights does TrinEval provide into the rote memorization and genuine capability learning in LLMs?
IV-A Experiment Setup
Models. We employ proprietary commercial LLMs for question reformulation in TrinEval, i.e., gpt-4o-2024-08-06(GPT) [13], qwen-max-2024-09-19(Qwen) [15, 14], deepseek-r1(Dpsk) [50] and qwen3-max-2026-01-23(Qwen3) [51]. For model evaluation, we utilize open-source LLMs to obtain token-level logits. Our experiments first focus on three widely adopted models: LlaMA2-7B (LlaMA) [7], Mistral-7B-v0.2 (Mistral) [11], and Vicuna-v1.5-7B (Vicuna) [12] due to their reported data contamination evidence [3, 7] on MMLU. Then, we further extend the results to the latest larger models, such as LlaMA3.1-8B [52], Qwen3-14B, and Qwen3-32B. These models are accessed from Hugging Face and implemented with the transformers library, allowing us to compute the log-probabilities of generated tokens for fine-grained analysis. All evaluations are conducted on a single NVIDIA A100 GPU with 40GB of memory. Default generation parameters are used, and greedy decoding is applied throughout to ensure reproducibility.
Benchmarks. We evaluate LLMs on the widely used MMLU benchmark [6], in light of growing evidence of data contamination [7] in many recent LLMs on this data. MMLU spans 57 subjects across STEM, humanities, social sciences, and others, providing a comprehensive assessment of model capabilities. We de-duplicate MCQs across subjects on MMLU. Also, MCQs from some subjects contain similar or identical options777e.g., the options of MCQs in the subject, moral_scenarios, are all identical (‘Wrong, Wrong’, ‘Wrong, Not wrong’, ‘Not wrong, Wrong’ and ‘Not wrong, Not wrong’).. With the provided 5-shot prompt, options of MCQs from these subjects can be easily extracted, leading to a high False-positive ratio. In order to avoid the influence of the few-shot prompt on the option extraction, we eliminate MCQs in which any of the options have appeared twice in the dataset, resulting in a test set of 14,006 unique questions. We also extend the evaluation on the widely used GSM8K benchmark [53], given the determinacy of its answers and the verified data contamination [3]. GSM8K consists of high-quality grade school math problems, focusing on multi-step arithmetic reasoning to rigorously evaluate the basic mathematical problem-solving ability of LLMs. In this way, we are able to evaluate the memorization of the OEQs according to the extraction of the official chain-of-thought(COT) sequences. We mainly select the test set of GSM8K as the evaluation benchmark. After filtering the duplicate OEQs, we obtain an OEQ set of 1,319 unique questions.
Evaluation. With commercial LLMs, we evaluate model performance by extracting predicted answers with regular expressions on MCQs. For open-source models, we access the output probability of the first generated token (i.e., corresponding to multiple-choice option IDs A/B/C/D) to compute quantitative performance metric for MCQs. For OEQs, we compute the perplexity of the answer string as the capability metric, only available with the open-source LLMs.
IV-B Q1. Is TrinEval Knowledge-preserving?
We begin by verify whether reformulation of TrinEval preserves essential knowledge, a prerequisite for reliable knowledge assessment. Specifically, our goal is to ensure that (1) the reformulation does not omit critical information that would cause originally correctly answered questions to be answered incorrectly, and (2) it does not introduce anomalous or unexpected content that could artificially inflate performance. We verified the above two objectives through both automatic analysis based on LLMs’ results and manual annotations, ensuring the robustness of TrinEval.
Q1.1 Automatic analysis
Figure 3: Knowledge-preserving validation of TrinEval. The green, blue, and overlapping regions represent the sets of MCQs correctly answered in the original format, TrinEval format, and both formats, respectively. Best viewed in color.
Following the complete TrinEval reformulation pipeline, we obtain 4,343 qualified MCQs with corresponding triplets using Qwen, and 4,645 MCQs with associated triplets using GPT. We then prompt both Qwen and GPT to answer their respective MCQs in both the original and reformulated triplet-based formats. Example MCQs and the questioning prompt templates in both formats are provided in Fig. 5 in Appendix B; results are summarized in Fig. 3. Here, we primarily employ robust proprietary LLMs to evaluate the knowledge retention capability of the TrinEval framework. This approach effectively mitigates potential interference from the models’ memorization effects, thereby ensuring accurate assessment of their intrinsic knowledge processing abilities and preventing erroneous conclusions that might arise from conflating memorization with genuine comprehension.
Among the qualified MCQs reformulated by Qwen and GPT, 4,133 and 4,409 are answered correctly in at least one of the two formats. Notably, the majority of these are correctly answered in both original and TrinEval formats, accounting for 90.3% with Qwen and 90.4% with GPT, indicating strong consistency in terms of problem-solving across formats. Only 121 questions for Qwen (2.9%) and 226 for GPT (5.1%) are answered correctly exclusively in the original format, suggesting the TrinEval reformulation does not omit essential information. Conversely, 278 questions for Qwen (6.7%) and 197 for GPT (4.5%) are answered correctly only in the reformulated format, which is comparable in general to the numbers for the original format alone. This further suggests that the reformulation does not introduce extra information that might artificially enhance model performance.
Q1.2 Manual annotation
Given the potential risk for the LLM-based Knowledge Preserving evaluation (line 4 and line 10 in 1), the proprietary LLMs might still be able to answer the MCQs without sufficient knowledge. Since this is still prompting LLMs who have been trained on these datasets and “know” the original content, human annotation is also applied. The annotation of each MCQ encompasses three subtasks: (1) answering the question in the re-organized form, (2) answering the question in the original form, and (3) verifying if the reformulation is Knowledge Preserving or not.
Metric Qwen GPT
TrinEval correct only 0.0% 0.0%
both correct 96.296% 96.667%
original correct only 3.704% 3.333%
K.P. score 4.101 4.369
TABLE I: Result of Human Annotation on LLM-based knowledge preserving (K.P.) evaluation in TrinEval
For efficient annotation, we implemented a stratified sampling procedure by selecting one MCQ per subject from all 56 MMLU subjects (as a temporary compromise for limited time, which will be expanded later) under both Qwen and GPT reorganization paradigms. This yielded 112 representative questions (2 systems × 56 subjects) for evaluation. Three human annotators independently performed dual-form assessments through: (1) Direct question answering with the reformed format first and the original format; (2) Knowledge Preservation (K.P.) scoring across two dimensions: i. Knowledge adequacy (sufficiency for accurate response), ii. removal of redundant content using a 5-point scale (1=unsatisfactory; 2=major information is missed or unnecessary information is incorporated, but part is still acceptable, 3=need to take some time to understand, but can still solve the MCQ, 4=an element properly belonging to one triplet component appears in another, but does not impact MCQ solving; 5=optimal). We use a continuous rather than a binary metric to mitigate the cognitive difference in the threshold between the annotators. Inter-rater reliability was ensured through consensus-building discussions prior to formal annotation. Final scores were aggregated using averaged values to further mitigate individual annotator bias. The results are shown in Tab. I below.
Our analysis reveals that over 95% of correctly answered MCQs maintained consistency across both original and paraphrased formats. Furthermore, human annotators rated our paraphrased questions mean K.P. score exceeding 4.0 (on a 5-point scale), which means that the reformulated MCQs only somehow influence the readability of humans but do not impact the solvability of the original format. This provides empirical validation that our proposed TrinEval methodology effectively preserves necessary knowledge elements from original MCQ formulations, while the influence of the LLM memorization during the evaluation is rather limited.
Experimental results under human annotation also reveal that Qwen underperforms GPT in key metrics, particularly in processing long-context texts, where it occasionally omits background information (evidenced by excessive ”N/A” assignments in the Context fields). This capability gap is further reflected in MCQ annotations: there are only MCQs that are merely correctly answered with the original format, except for the correct MCQs with both formats. Collectively, these findings provide strong evidence that TrinEval preserves the core knowledge necessary for answering, satisfying the requirements for reliable capability evaluation.
IV-C Q2. Can TrinEval Reduce Memorization?
((a))
((b))
((c))
((d))
((e))
((f))
Figure 4: Results of memorization evocation (evoc.) under dataset-related contexts, with green and blue curves referring to the memorization difference $\Delta F_{m}$ in the original and TrinEval formats, respectively. In the x-axis, ‘clean’, ‘meta’, ‘dev-fsp’, and ‘seq-fsp’ stand for without dataset-related context, with the name of the dataset, with few-shot prompt from the development set, and with few-shot prompt from the test set ahead of the testing question. These results of $\Delta F_{m}$ indicate the growing memorization effect given the increasing dataset-related information in general. However, the $\Delta F_{m}$ by TrinEval under the strongest memory evocation context remains consistently lower than the one in the original format, e.g., the red dashed line.
In this subsection, we investigate whether the proposed TrinEval reformulation can suppress unnecessary memorization, thereby isolating and revealing the LLMs’ genuine capabilities under various circumstances. Following Bordt et al. [4], we introduce dataset-specific cues into the questioning prompts and assess to what extent TrinEval mitigates memorization elicited by such prompts.
We suppose that unintentional data contamination arises from crawling dataset pages (e.g., Hugging Face) during the compilation of LLM pretraining datasets. When researchers are organizing the pretraining corpus, one or more neighboring original data samples would be truncated and concatenated sequentially into a pretraining sample. Thus, according to Carlini’s theory [1] and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
Thus, following the previous study [4], to evoke memorization, we embed dataset-related metadata into the input prompts, i.e., the dataset name and in-context few-shot examples drawn from the same dataset, to find out if the reformulation of the TrinEval framework can eliminate the memorization effect. For this set of experiments, we focus on LlaMA, Mistral, and Vicuna, as there exists verified evidence of data contamination and open-source implementations allow access to token-level output probabilities, enabling us to compute the memorization metric $F_{m}$. Since $F_{m}$ lacks a defined absolute zero point indicating complete absence of memorization, we take the $F_{m}$ value of the original MCQ format without any additional context (i.e.,, vanilla MCQ) as a reference baseline. We visualize the average change in $F_{m}$ for each memorization-evoking method relative to this baseline.
Specifically, we hypothesize that unintentional data contamination may occur when LLMs are pretrained on datasets scraped from public sources (e.g.,, Hugging Face), where nearby examples from the same dataset may be concatenated during corpus construction. Following the pretraining mechanism of next-token prediction and prior findings from Carlini et al. [1], we posit that the inclusion of dataset-similar samples in the prompt may inadvertently trigger memorization. Accordingly, we design a sequence of memorization-evoking perturbations by progressively increasing the contextual cues: (1) providing only the dataset name, (2) adding few-shot examples from the same dataset (i.e.,, samples in the development set or ahead examples in the test set of the same subject of MMLU), and (3) combining both types of context. For each MCQ, we compute the change in $F_{m}$ relative to the reference baseline, capturing the degree to which memorization is elicited by specific context in original or TrinEval formats. Results are illustrated in Fig. 4.
Consistent with the observations by Bordt et al. [4], results show that $F_{m}$ increases as more dataset-specific context is introduced, indicating stronger memorization effects. However, across all three tested LLMs, the $\Delta F_{m}$ curve for TrinEval remains consistently lower than that of the original MCQ format (i.e., the red dashed line) regardless of various context. Notably, even under the strongest memorization-evoking setting, TrinEval ’s absolute $F_{m}$ still stays below the vanilla baseline. These findings provide compelling evidence that TrinEval significantly mitigates the influence of memorization, effectively disentangling recall from genuine model capability.
IV-D Q3. TrinEval’s Findings on LLM Memorization and Genuine Capability Learning
((a))
((b))
((c))
((d))
((e))
((f))
Figure 5: The distribution of MCQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$. According to the values of $F_{m}$ and $F_{c}$, we separate MCQs into 25 groups and visualize the distribution from weak to strong with heatmaps. Same as below.
In this subsection, we aim to examine the relationship between memorization and genuine capability in LLMs, using the metrics $F_{m}$ (memorization) and $F_{c}$ (capability).
Q3.1 TrinEval’s findings on MCQs
Since commercial proprietary models do not expose full vocabulary-level output probabilities, we focus on open-source LLMs for computing these metrics. After computing the $F_{m}$ and $F_{c}$ scores for all qualified MCQs in TrinEval format, we divide the values of $F_{m}$ and $F_{c}$ into five equal intervals, respectively, resulting in 25 distinct groups of MCQ samples based on their joint distribution. We then use heatmaps to visualize the distribution of samples across these groups, revealing the relationship between memorization and capability for each tested model.
As shown in Fig.5, the majority of MCQs cluster in the lower-left and upper-right corners of the heatmap. For instance, in the results of LLaMA using Qwen-extracted triplets, 38.57% of MCQs fall within the combined lower-left and upper-right $2\times 2$ grid regions, yielding a Pearson correlation of -0.7755 (p-value <0.05<0.05) for $F_{m}$ vs. $F_{c}$ of examples within the two square regions; expanding to the $3\times 3$ regions increases coverage to 74.17%, with a stronger correlation of -0.8124 (p-value <0.05<0.05). Similarly, for Mistral on Qwen triplets, 44.90% of MCQs lie in the two $2\times 2$ corner regions with a correlation of -0.8722 (p-value <0.05<0.05), and 80.82% in the two $3\times 3$ corner regions with a correlation of -0.8794 (p-value <0.05<0.05).
((a))
((b))
((c))
((d))
((e))
((f))
Figure 6: The distribution of MCQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$ with latest larger models.
We analyze the distribution in detail by revealing the ratio of MCQs within the upper right and lower left 2 $\times$ 2 and 3 $\times$ 3 squares as well as the Pearson correlations between the $F_{m}$ and $F_{c}$ of these MCQs. We take the MCQs within the lower-left 2 $\times$ 2 squares in Fig. 5 as the ones memorized by rote and compute the ratio of 20.5% by averaging the percentage of all evaluated LLMs. Similarly, 19.6% is computed as the ratio of MCQs within the upper-right 2 $\times$ 2 squares with genuine capability. All of these MCQs also exhibit a tendency towards a negative correlation between the capabilities and memorization of LLMs (all Pearson correlation values are computed with p-values at the 0.05 level). Additional statistics results are reported in Tab. III in Appendix C. Further, we also plot the curve of the separation of the rote memorization and genuine capability learning MCQs in Appendix C. With each progressing $F_{m}$ as threshhold, the tested MCQs exhibit prominent $F_{c}$ gaps between the two groups, showing significant difference in learning extremes.
These findings consistently indicate a negative correlation between memorization and capability: MCQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance, suggesting overfitting or shallow recall. Still, we would argue that knowledge acquisition in LLMs is likely a continuous spectrum rather than a strict binary. Our preliminary investigation also highlighted the limitations of a discrete classification, and thus, we apply the continuous quantification metric $F_{m}$ and $F_{c}$. As seen in the heatmaps, some questions fall in intermediate regions, indicating varying degrees of memorization and understanding. Thus, while we focus on the two extremes for analytical clarity, we agree that other intermediate or hybrid categories may exist.
Q3.2 Can TrinEval’s findings be extended to latest larger LLMs?
Based on previous research [7, 3], the aforementioned experiments were conducted in a relatively deterministic experimental environment, i.e., confirmed data contamination evidence. Therefore, we attempt to extend the above results to the latest larger LLMs. Here, we mainly reformulate the MMLU MCQs based on Dpsk and Qwen3, obtaining 6,115 and 5,905 MCQs, respectively, and verify the results based on LlaMA3.1-8B, Qwen3-14B, and Qwen3-32B. The experimental procedure remains identical to that described in the preceding experiment. Results are shown in Fig. 6. We can see that the majority of MCQs also cluster in the lower-left and upper-right corners of the heatmap, which is evident and consistent across our tests. For instance, in the results of LlaMA3.1 using Qwen3-extracted triplets, 38.34% of the MCQs fall within the two opposite corner regions of the $2\times 2$ grid. For the examples in these two regions, the Pearson correlation between $F_{m}$ and $F_{c}$ is -0.7771 (p-value <0.05<0.05). Extending the analysis to the $3\times 3$ corner regions increases coverage to 78.34%, yielding a correlation of 0.5506 (p-value <0.05<0.05). Similarly, for Qwen-14B on Qwen triplets, 39.04% of the MCQs lie in the two $2\times 2$ corner regions, with a correlation of -0.7260 (p-value <0.05<0.05); and 79.02% fall in the two $3\times 3$ corner regions, with a correlation of -0.4905 (p-value <0.05<0.05).
((a))
((b))
((c))
((d))
((e))
((f))
Figure 7: The distribution of OEQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 8: The distribution of OEQs based on memorization metric $F_{m}$ vs. capability metric $F_{c}$ with latest larger models.
Q3.3 Can TrinEval’s findings be extended to OEQs?
Subsequently, we pose a new question: can the above findings on MCQs be extended to OEQs? To answer this question, we apply TrinEval on the GSM8K dataset. For $F_{m}$, instead of the original option content in MCQs, we select the bottom K% token probabilities of the official CoT seuqence for each OEQ to evluate the memorization of LLMs. As to $F_{c}$, given that GSM8K is a mathematics problem benchmark with a definite answer, we utilize the perplexity of the answer sequence. For open-ended questions (OEQs) that admit multiple correct answer formulations and require semantic correctness evaluation, we believe that our TrinEval can also be readily adapted to corresponding measurement. We utilize TrinEval to reformulate the OEQs and obtain 1282 qualified questions and use them to evaluate LlaMA, Mistral, Vicuna, as well as latest larger models, Llama3.1-8B, Qwen3-14B, and Qwen3-32B. The results are shown in Fig. 7 and Fig. 8.
We surprisingly find that on OEQs, most LLMs still exhibit the same phenomena, i.e., OEQs with lower memorization scores are more likely to reflect genuine problem-solving capability, whereas high memorization levels correspond to degraded performance. Among them, for LLaMA on the triplet questions extracted by Dpsk, 18.04% of the questions exhibited rote memorization, while 17.62% were genuinely mastered. For Qwen3‑32B on the same Dpsk‑extracted triplets, 17.62% of the questions showed rote memorization, and 17.53% were genuinely mastered. However, for some results, for instance, LLaMA3.1‑8B on the triplet extracted by Qwen3, only 15.76% of the questions were classified as rote memorization, falling below the 16% baseline in the $5\times 5$ heatmap. We attribute this to the following reasons. First, comparing Fig. 7 and Fig. 8, we observe that the distinction between rote memorization and genuine capability learning is less pronounced in the latest larger LLMs than in other LLMs. This may be because, as suggested by recent research [55, 56, 21], LLMs are increasingly designed to be trained on reformulated corpus during pretraining, thereby reducing the formation of such rote memorization. Second, Several studies [54] have also found that overly formatted corpora can lead to poorer generalization in LLMs. Thus, compared to MCQs, OEQs are less structured and thus may be less susceptible to rote memorization. Nevertheless, the conclusions drawn from most LLMs remain consistent with our earlier experimental findings on MCQs, suggesting that LLMs also exhibit rote memorization in the context of OEQs.
Q3.4 How to interpret TrinEval’s findings?
To interpret these results, we draw an analogy between LLM behavior and the human memory system, which comprises two key components: Short-Term Memory (STM) and Long-Term Memory (LTM) [46]. Neurobiological studies reveal that STM relies on transient synaptic protein synthesis with limited temporal persistence and functional scalability. In contrast, LTM is constructed through stabilized neuronal memory traces that constitute an enduring knowledge framework. This neural architecture not only supports STM operations as a cognitive substrate but also enables sophisticated information generalization across diverse contexts. This dichotomy aligns with recent observations in LLMs. As shown by Allen et al. [18] and Ovadia et al. [21], models trained on corpora with diverse rephrasings exhibit stronger generalization than those trained solely on original formulations. When trained on a single fixed corpus format, LLMs, due to their strong memorization capacity, tend to activate their STM system to memorize at the token level, capturing surface patterns rather than abstract knowledge. In this sense, LLMs are potential rote learners. Reformulations like TrinEval appear to facilitate the encoding of knowledge in a more principled, structured and reusable form, akin to LTM, thereby supporting more generalized and robust evaluation.
Figure 9: Averaged distance of each MCQs between the closest $1\%$ MCQs’ embeddings. ‘Rote Memorization’ refers to MCQ within lower left $2\times 2$ squares that typically exhibits a high memorization metric $F_{m}$ and a low capability $F_{c}$ while ‘Genuine Capability Learning’ stands for MCQ lies within the upper right $2\times 2$ squares with low $F_{m}$ but high $F_{c}$.
To further validate our hypothesis, we analyze the semantic proximity of MCQs within the qualified MMLU dataset by computing their embeddings and measuring the average distance to their closest 1% samples. Our underlying assumption is that LLMs could better master a knowledge point if it is expressed in various formats in the training corpus. As a result, these knowledge ‘rephrases’ are also encoded densely via LLMs. As there are 57 different subjects within the MMLU dataset, we argue that unrelated sequences would lead to increased embedding distance even though they are among the mastered knowledge points. Here, we try to filter out the unrelated samples for each MCQ and thus keep the closest samples at the 1e-2 level in order to make sure there are not too many unrelated samples incorporated. To highlight this result more prominently, we employ a relatively stringent filtering strategy and made it 1% of the closest samples, aiming to retain the samples that are semantically correlated in knowledge description. We then compare the mean embedding distances of MCQs located in the lower-left (strong memorization, weak capability) and upper-right (weak memorization, strong capability) $2\times 2$ regions in Fig. 5. The corresponding results are presented in Fig. 9, where RM stands for rote memorization, and GCL stands for genuine capability learning. In order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5% in Tab. II in Appendix D. Interestingly, we observe that the average embedding distance among the Genuine Capability Learning MCQs is much lower than (nearly half) that of the Rote Memorization MCQs. This suggests that memorized MCQs are more sparsely distributed in the embedding space, while non-memorized, capability-driven MCQs tend to form tighter semantic clusters. The observation also aligns well with the cognitive analogy of STM and LTM: rote memorization leads to fragmented, context-specific encoding, whereas genuine capability emerges from structured, reusable representations. Also, we can see that the more samples we incorporated, the higher the average distance of the closest embeddings grows. Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings. As a conclusion, though it is widely believed that memorization may lead to better but cheating performance of LLMs, we find that the more LLMs memorize, the worse they are at solving problems.
V Conclusion
[…]